OpenStudy (anonymous):

the ratio of the areas of two similar polygons is 49:16 if the perimeter of the first polygon is 22cm what is the perimeter of the second polygon

OpenStudy (anonymous):

49/16 = A1/A2 sqrt(49) / sqrt(16) = P1/P2 = 7/4 = 22/x So now, just solve for "x" 7x = (4)(22) x = (4)(22)/(7)

OpenStudy (anonymous):

its 12.57

OpenStudy (anonymous):

You got it! Good job!

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

uw!